1 kg of a platinum-iridium has 39 mm in height and 39 mm in diameter. The

volume of cylinder is V = πr2h. Calculate the density of the cylinder in SI unit.


1
Expert's answer
2021-08-02T07:14:05-0400

Calculate the density of the cylinder

"\rho=\frac{m}{V}=\frac{m}{\pi r^2h}\\
=\frac{1\:\rm kg}{3.14*(0.039/2\:\rm m)^2*(0.039\:\rm m)}=21475\:\rm kg/m^3"